OpenStudy (anonymous):

Two factory plants are making TV panels. Plant A produced 2000 panels. Two percent of the panels from plant A and 5% of the panels from plant B were defective.How many panels did plant B produce if the over all percentage of defective panels from both plants was 4%?

OpenStudy (akashdeepdeb):

Let number of panels of Plant B = x !! :D So how many are defective from plant A? = 2 * 2000/100 = 40 panels!! :D And how many are defective from Plant B = 5x/100 Now they are saying that the total 'PERCENTAGE' of defective panels from both is 4% So 4(2000 + x)/100 = 5x/100 + 40!! :D or 8000 + 4x = 5x + 4000 x = 4000!! So the number of panels produced by Plant B = 4000 panels!! :D Getting me? :)